OpenStudy (mven):

Find the area of a circle whose area equals the half of the sum of the areas of two circles with radii 2 and 3. 13 6.5 5

OpenStudy (michele_laino):

hint: area of circle with radius 2 is: \[3.14 \times 2 \times 2 = ...?\]

OpenStudy (mven):

12.56

OpenStudy (mven):

thats what i did last time got the same think but is not one of the answeres

OpenStudy (mven):

@Michele_Laino

OpenStudy (mven):

Would it be 6.27pi?

OpenStudy (mven):

6.28pi not 6.27

OpenStudy (michele_laino):

ok! Now area of circle with radius 3, is: \[3.14 \times 3 \times 3 = ...?\]

OpenStudy (mven):

28.26

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now the half of the sum of those areas is: \[\frac{{28.26 + 12.56}}{2} = ...?\]

OpenStudy (mven):

20.41

OpenStudy (michele_laino):

I think that is your answer

OpenStudy (mven):

the answers is is having me chose from are 13pi 6pi and 5pi

OpenStudy (michele_laino):

it is simple: what is \[\frac{{20.41}}{{3.14}} = ...?\]

OpenStudy (mven):

ahh 6.5 which would be 6.5 pi?

OpenStudy (michele_laino):

yes!

OpenStudy (mven):

awesome thank you

OpenStudy (michele_laino):

thank you!